OpenStudy (anonymous):

Hey guys i got a question, if you could help me that would be great! You buy a pot and its lid for a total of $ 11. The sales person tells you that the pot by itself costs $ 10 more than the lid. The price of the pot is $ ? and the price of the lid is ?

OpenStudy (anonymous):

Pot = 10$ Lid = 1$

OpenStudy (anonymous):

Oh, Wait

OpenStudy (anonymous):

i tried that...

OpenStudy (anonymous):

Doesn't seam possible

OpenStudy (anonymous):

is there an equation that we can use to get a different answer

OpenStudy (anonymous):

Ut oh amistre Wrighting

OpenStudy (amistre64):

P + L = 11 P = L + 10 (L+10)+L = 11 2L + 10 = 11 2L = 11-10 2L = 1 Lid = 1 dollar, Pot = 10 dollars....

OpenStudy (amistre64):

make that 50 cents and 11.50 lol

OpenStudy (amistre64):

10.50.....ack!!!

OpenStudy (anonymous):

If Lid = 1$ then Pot has to be 11$ = 12$ total but he bought it for 11$

OpenStudy (amistre64):

lid = 0.50; pot = 10.50

OpenStudy (anonymous):

Ahhh

OpenStudy (amistre64):

it works out on paper, but when I try to type ity in it gets all discumbobulated :)

OpenStudy (anonymous):

ok thanks guys, amistre so how did you get that, it says its right i just want to know so i can figure it out

OpenStudy (amistre64):

Well, I used the information given to construct a logical model. We know the price of the pot(P) and lid(L) added together is $11. P + L = $11 We know that the P is equal to the lid(L) + $10; so replace P with L + 10 (L + 10) + L = 11 Combine like terms ; L + L = 2L 2L + 10 = 11 subtract (10) from both sides to keep everything in balance and to get the Ls by them self. 2L +10-10 = 11-10 2L = 1 to find what L equals, we divide each side by (2) to keep them balanced. (2L)/2 = 1/2 L = 1/2 of a dollar. I know that .50 is exactly one-half of a dollar :) Since I know the price of a lid(L) I can now add $10 to it to determine the price of the pot(P). P = .50 + 10.00 = 10.50.